# The Exchangeability Assumption for Permutation Tests of Multiple Regression Models

### v1

@@ 6 Acknowledgements
This work was performed in part using high-performance computing equipment at Amherst College obtained under National Science Foundation Grant No. 2117377. We thank Shiya Cao and Lindsay Poirier for pointing us toward the bridges data. We thank Shiya Cao for useful comments on an earlier draft of the manuscript.
@@ 8 Reproducibility statement
The code used to undertake the Bridges permutation analysis in Section 3 has been made available at https://osf.io/7zcfu/.
Johanna Hardin: https://orcid.org/$0$-0001-6251-1955
Nicholas Horton: https://orcid.org/$0$-0003-3332-4311
A Appendix
Claim. The correlation between $Y^{*}$ and $X_{2}$ is dependent on the correlation between $X_{1}$ and $X_{2}$.
@@ A Appendix
$\rho(Y^{*},X_{2})$ $=\frac{cov(Y^{*},X_{2})}{\sqrt{var(Y^{*})var(X_{2})}}$
Breaking down each part of the correlation between $Y^{*}$ and $X_{2}$, we compute both $cov(Y^{*},X_{2})$ and $var(Y^{*}).$
$cov(Y^{*},X_{2})$ $=cov(b_{0\cdot 1}+b_{1}X_{1}+R^{*}_{Y\cdot 1},X_{2})$
$=cov(b_{0\cdot 1},X_{2})+cov(b_{1}X_{1},X_{2})+cov(R^{*}_{Y|1},X_{2})$
$\approx cov(b_{1}X_{1},X_{2})$ (4)
$\approx b_{1}\cdot cov(X_{1},X_{2})$ (5)
$=b_{1}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})var(X_{2})}$
Note that $b_{0}\cdot 1$ and $b_{1}$ are a random variables because they are a statistics, so Equations (5,6,7) are all approximate. In Equation (4), we assume that $b_{0\cdot 1}$ and $R^{*}_{Y\cdot 1}$ are both independent of $X_{2}$, which leads to $cov(b_{0\cdot 1},X_{2})=cov(R^{*}_{Y\cdot 1},X_{2})=0$.
$var(Y^{*})$ $=var(b_{0\cdot 1}+b_{1}X_{1}+R^{*}_{Y|1})$
$\approx(b_{1})^{2}var(X_{1})+var(R^{*}_{Y|1})$ (6)
$var(R^{*}_{Y\cdot 1})$ is broken down into pieces.
$var(R^{*}_{Y\cdot 1})$ $=var(R_{Y\cdot 1})$
$=var(Y-b_{0\cdot 1}-b_{1}X_{1})$
$\approx var(Y)+(b_{1})^{2}var(X_{1})-2b_{1}cov(Y,X_{1})$ (7)
$cov(Y,X_{1})$ $=cov(\beta_{0\cdot 1,2}+\beta_{1\cdot 2}X_{1}+\beta_{2\cdot 1}X_{2}+\varepsilon,X_{1})$
@@
$=\beta_{1\cdot 2}var(X_{1})+\beta_{2\cdot 1}cov(X_{2},X_{1})$ (8)
Plugging $cov(Y,X_{1})$ into $var(R^{*}_{Y\cdot 1})$ and $var(R^{*}_{Y\cdot 1})$ into $var(Y^{*})$, we can approximate $var(Y^{*}).$
@@
$=2b_{1}(b_{1}-\beta_{1\cdot 2})var(X_{1})+var(Y)-2b_{1}\beta_{2\cdot 1}cov(X_{2},X_{1})$
$=2b_{1}[(b_{1}-\beta_{1\cdot 2})var(X_{1})-\beta_{2\cdot 1}cov(X_{1},X_{2})]+var(Y)$
@@
$\rho(Y^{*},X_{2})$ $=\frac{cov(Y^{*},X_{2})}{\sqrt{var(Y^{*})var(X_{2})}}$
$\approx\frac{b_{1}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})var(X_{2})}}{\sqrt{var(Y^{*})var(X_{2})}}$
$=\frac{b_{1}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})}}{\sqrt{var(Y^{*})}}$
$=\frac{b_{1}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})}}{\sqrt{2b_{1}[(b_{1}-\beta_{1\cdot 2})var(X_{1})-\beta_{2\cdot 1}cov(X_{1},X_{2})]+var(Y)}}.$ (9)

### v2

@@ 5 Acknowledgements
This work was performed in part using high-performance computing equipment at Amherst College obtained under National Science Foundation Grant No. 2117377. We thank Shiya Cao and Lindsay Poirier for pointing us toward the bridges data and Shiya Cao for useful comments on an earlier draft of the manuscript and the reviewers and associate editor for helpful suggestions.
@@ 7 Reproducibility statement
The code used to undertake the simulation study reported in Table LABEL:tbl-simresults and the supplementary resources (bridges analysis) has been made available at https://osf.io/7zcfu/?view.
Appendix 0 Appendix
Here we review different approaches for carrying out permutation tests for multiple linear regression models.
1 Permutation algorithms
Permute $Y$
Summary of algorithm when permuting $Y$ (Manly)
  1. 1.
    Fit the original model and obtain coefficient estimates ($b_{0\cdot 1,2}$, $b_{1\cdot 2}$, and $b_{2\cdot 1}$) and corresponding standard error estimates ($SE(b_{0\cdot 1,2})$, $SE(b_{1\cdot 2})$, and $SE(b_{2\cdot 1})$):
    $\widehat{Y}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}$
  2. 2.
    Permute $Y$ to obtain $Y^{*}$.
  3. 3.
    Fit a model on the permuted $Y^{*}$ values to obtain permuted coefficient estimates ($b^{*}_{0\cdot 1,2}$, $b^{*}_{1\cdot 2}$, and $b^{*}_{2\cdot 1}$) and corresponding standard error estimates ($SE(b^{*}_{0\cdot 1,2})$, $SE(b^{*}_{1\cdot 2})$, and $SE(b^{*}_{2\cdot 1})$):
    $\widehat{Y}^{*}=b^{*}_{0\cdot 1,2}+b^{*}_{1\cdot 2}X_{1}+b^{*}_{2\cdot 1}X_{2}$
  4. 4.
    Repeat steps 2 and 3 $P$ times. For example, $P$ = 1000.
  5. 5.
    From the $P$ copies of $b^{*}_{1\cdot 2}$ and $P$ copies of $SE(b^{*}_{1\cdot 2})$, calculate $P$ copies of $t^{*}$ to form the permuted null sampling distribution:
    $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$ (1)
  6. 6.
    Compare the observed test statistic to the permuted null sampling distribution from step 5:
    $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$
Permute $X_{1}$
Summary of algorithm when permuting $X_{1}$ (Draper and Stoneman)
  1. 1.
    Fit the original model and obtain coefficient estimates and corresponding standard error estimates:
    $\widehat{Y}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}$
  2. 2.
    Permute $X_{1}$ to obtain $X^{*}_{1}$.
  3. 3.
    Fit a model on the permuted $X_{1}^{*}$ values to obtain permuted coefficient estimates and corresponding standard error estimates:
    $\widehat{Y}=b^{*}_{0\cdot 1,2}+b^{*}_{1\cdot 2}X^{*}_{1}+b^{*}_{2\cdot 1}X_{2}$
  4. 4.
    Repeat steps 2 and 3 $P$ times.
  5. 5.
    From the $P$ copies of $b_{1\cdot 2}^{*}$ and $P$ copies of $SE(b_{1\cdot 2}^{*})$, calculate $P$ copies of $t^{*}$ to form the permuted null sampling distribution:
    $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$ (2)
  6. 6.
    Compare the observed test statistic to the permuted null sampling distribution from step 5:
    $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$
Permute reduced model residuals
Summary of algorithm when permuting reduced model residuals (Freedman and Lane)
  1. 1.
    Fit the original model on $X_{2}$ only and obtain coefficient estimates and corresponding standard error estimates of the reduced model:
    $\widehat{Y}=b_{0\cdot 2}+b_{2}X_{2}$
  2. 2.
    Let the residuals $R_{Y\cdot 2}=Y-b_{0\cdot 2}-b_{2}X_{2}$, and permute $R_{Y\cdot 2}$ to obtain $R^{*}_{Y\cdot 2}.$ Define the permuted outcome variable as $Y^{*}=b_{0\cdot 2}+b_{2}X_{2}+R^{*}_{Y\cdot 2}.$
  3. 3.
    Fit a model on the permuted $Y^{*}$ values to obtain permuted coefficient estimates and corresponding standard error estimates:
    $\widehat{Y}^{*}=b^{*}_{0\cdot 1,2}+b^{*}_{1\cdot 2}X_{1}+b^{*}_{2\cdot 1}X_{2}$
  4. 4.
    Repeat steps 2 and 3 $P$ times.
  5. 5.
    From the $P$ copies of $b_{1\cdot 2}^{*}$ and $P$ copies of $SE(b_{1\cdot 2}^{*})$, calculate $P$ copies of $t^{*}$ to form the permuted null sampling distribution:
    $t^{*}=\frac{b_{1\cdot 2}^{*}-0}{SE(b_{1\cdot 2}^{*})}$ (3)
  6. 6.
    Compare the observed test statistic to the permuted null sampling distribution from step 5:
    $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$
Claim. The correlation between $Y^{*}$ and $X_{1}$ is dependent on the correlation between $X_{1}$ and $X_{2}$.
  1. (1)
    The permutation scheme is designed to create new values for the coefficient on $X_{1}$ under the setting that $X_{1}$ and $Y$ are uncorrelated given $X_{2}$. Repeated sets of the “permuted” $Y^{*}$ values are generated by adding noise (permuted residuals) to the fitted values from the model on $X_{2}$ only. At first glance, it seems as though the method creates $Y^{*}$ values that are associated with $X_{2}$ (because they built from the $X_{2}$-model fitted values) and not associated with $X_{1}$ (because they are built from a model that completely ignores $X_{1}$). However, if $X_{1}$ and $X_{2}$ are correlated, then $Y^{*}$ values that are correlated with $X_{2}$ will naturally be correlated with $X_{1}$.
  2. (2)
    In Appendix 2, we give a proof sketch outlining the derivation for the covariance between $Y^{*}$ and $X_{1}$. Although the proof is only outlined, the dependence on the correlation between $X_{1}$ and $X_{2}$ is clear. The approximate value of $\rho(Y^{*},X_{1})$ derived in Appendix 2 is validated empirically (results not shown); see Ye (2023) for empirical results.
Permute full model residuals
Summary of algorithm when permuting full model residuals (ter Braak)
  1. 1.
    Fit the original model and obtain coefficient estimates and corresponding standard error estimates of the full model:
    $\widehat{Y}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}$
  2. 2.
    Let the residuals $R_{Y\cdot 1,2}=Y-b_{0\cdot 1,2}-b_{1\cdot 2}X_{1}-b_{2\cdot 1}X_{2}$, and permute $R_{Y\cdot 1,2}$ to obtain $R^{*}_{Y\cdot 1,2}.$ Define the permuted outcome variable as $Y^{*}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}+R^{*}_{Y\cdot 1,2}.$
  3. 3.
    Fit a model on the permuted $Y^{*}$ values to obtain permuted coefficient estimates and corresponding standard error estimates:
    $\widehat{Y}^{*}=b^{*}_{0\cdot 1,2}+b^{*}_{1\cdot 2}X_{1}+b^{*}_{2\cdot 1}X_{2}$
  4. 4.
    Repeat steps 2 and 3 $P$ times.
  5. 5.
    From the $P$ copies of $b_{1\cdot 2}^{*}$ and $P$ copies of $SE(b_{1\cdot 2}^{*})$, calculate $P$ copies of $t^{*}$ to form the permuted null sampling distribution:
    $t^{*}=\frac{b_{1\cdot 2}^{*}-b_{1\cdot 2}}{SE(b_{1\cdot 2}^{*})}$ (4)
  6. 6.
    Compare the observed test statistic to the permuted null sampling distribution from step 5:
    $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$
The full model residual method has a bootstrap flavor, but the permutation of the residuals is done without replacement (whereas bootstrapping is done with replacement). Similar to the development of other bootstrapping methods (Efron and Tibshirani 1994, 87), the underlying technical condition for ter Braak’s method is that:
$F(x)_{\big{(}\frac{b_{1\cdot 2}^{*}-b_{1\cdot 2}}{SE(b_{1\cdot 2}^{*})}\big{)}}\approx F(x)_{\big{(}\frac{b_{1\cdot 2}-\beta_{1\cdot 2}}{SE(b_{1\cdot 2})}\big{)}},$ (5)
where $F$ is the cumulative distribution function of the random variable described in the subscript of the function.
The sampling distribution of $t^{*}=\frac{b^{*}_{1\cdot 2}-b_{1\cdot 2}}{SE(b_{1\cdot 2}^{*})}$ approximates the sampling distribution of $t=\frac{b_{1\cdot 2}-\beta_{1\cdot 2}}{SE(b_{1\cdot 2})}$ due to the theoretical underpinnings from bootstrapping. That is, the variability of $b_{1\cdot 2}^{*}$ around $b_{1\cdot 2}$ mimics the variability of $b_{1\cdot 2}$ around $\beta_{1\cdot 2}.$
Under Equation (5), we can use the $t^{*}$ distribution constructed from many permutations of the same dataset to carry out a hypothesis test. If $H_{0}\!\!:\beta_{1\cdot 2}=0$ is true, then $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$ would be a likely value in the $t^{*}$ distribution, corresponding to a non-significant $p$-value, resulting in a failure to reject $H_{0}.$
However, if $H_{A}\!\!:\beta_{1\cdot 2}\neq 0$ is true and, say, $\beta_{1\cdot 2}=47,$ then we would expect $\hat{t}=\frac{b_{1\cdot 2}-47}{SE(b_{1\cdot 2})}$ to lie well within the $t^{*}$ distribution, while $t_{obs}=\frac{b_{1\cdot 2}-0}{SE(b_{1\cdot 2})}$ would lie on the margins, leading to a small $p$-value that concludes the test by rejecting $H_{0}.$
Furthermore, ter Braak (1990) synthesizes equations from various authors – Efron (1982) for the bootstrap, along with Cox and Hinkley (1974) and Lehmann and D’Abrera (1975) for the permutation – to make the following statements regarding the expected value and variance of the estimated slope coefficients in the bootstrapping versus permutation settings. $b_{1\cdot 2}^{+}$ corresponds to the coefficient derived from a standard bootstrap (full model), while $b_{1\cdot 2}^{*}$ corresponds to the permutation of the residuals under the full model.
$E^{+}(b_{1\cdot 2}^{+})$ $=E^{*}(b_{1\cdot 2}^{*})=b_{1\cdot 2}$ (6)
$var^{+}(b_{1\cdot 2}^{+})$ $=(1-1/n)var^{*}(b_{1\cdot 2}^{*}).$ (7)
Equation (6) indicates that the expected values of $b_{1\cdot 2}^{+}$ and $b_{1\cdot 2}^{*}$ are both $b_{1\cdot 2}$. Equation (7) suggests that the variance of $b_{1\cdot 2}^{+},$ the bootstrapped estimate, is smaller than the variance of $b_{1\cdot 2}^{*}.$ Hence, ter Braak uses Equations (6) and (7), along with the order property that $b_{1\cdot 2}^{+}$ and $b_{1\cdot 2}^{*}$ differ by $O(1/n)$ in second or higher order moments, to justify his proposal of the full model residual permutation strategy.
2 Impact of correlation between $X_{1}$ and $X_{2}$
Claim. The correlation between $Y^{*}$ and $X_{1}$ is dependent on the correlation between $X_{1}$ and $X_{2}$.
@@ 2 Impact of correlation between $X_{1}$ and $X_{2}$
$\rho(Y^{*},X_{1})$ $=\frac{cov(Y^{*},X_{1})}{\sqrt{var(Y^{*})var(X_{1})}}$
Breaking down each part of the correlation between $Y^{*}$ and $X_{1}$, we compute both $cov(Y^{*},X_{1})$ and $var(Y^{*}).$
$cov(Y^{*},X_{1})$ $=cov(b_{0\cdot 2}+b_{2}X_{2}+R^{*}_{Y\cdot 2},X_{1})$
$=cov(b_{0\cdot 2},X_{1})+cov(b_{2}X_{2},X_{1})+cov(R^{*}_{Y\cdot 2},X_{1})$
$\approx cov(b_{2}X_{2},X_{1})$ (8)
$\approx b_{2}\cdot cov(X_{1},X_{2})$ (9)
$=b_{2}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})var(X_{2})}$
Note that $b_{0\cdot 2}$ and $b_{2}$ are random variables because they are statistics, so Equations (9,10,11) are all approximate. In Equation (8), we assume that $b_{0\cdot 2}$ and $R^{*}_{Y\cdot 2}$ are both independent of $X_{1}$, which leads to $cov(b_{0\cdot 2},X_{1})=cov(R^{*}_{Y\cdot 2},X_{1})=0$.
$var(Y^{*})$ $=var(b_{0\cdot 2}+b_{2}X_{2}+R^{*}_{Y\cdot 2})$
$\approx(b_{2})^{2}var(X_{2})+var(R^{*}_{Y\cdot 2})$ (10)
$var(R^{*}_{Y\cdot 2})$ is broken down into pieces.
$var(R^{*}_{Y\cdot 2})$ $=var(R_{Y\cdot 2})$
$=var(Y-b_{0\cdot 2}-b_{2}X_{2})$
$\approx var(Y)+(b_{2})^{2}var(X_{2})-2b_{2}cov(Y,X_{2})$ (11)
$cov(Y,X_{2})$ $=cov(\beta_{0\cdot 1,2}+\beta_{1\cdot 2}X_{1}+\beta_{2\cdot 1}X_{2}+\varepsilon,X_{2})$
@@
$=\beta_{2\cdot 1}var(X_{2})+\beta_{1\cdot 2}cov(X_{1},X_{2})$ (12)
Plugging $cov(Y,X_{2})$ into $var(R^{*}_{Y\cdot 2})$ and $var(R^{*}_{Y\cdot 2})$ into $var(Y^{*})$, we can approximate $var(Y^{*}).$
@@
$=2b_{2}(b_{2}-\beta_{2\cdot 1})var(X_{2})+var(Y)-2b_{2}\beta_{1\cdot 2}cov(X_{1},X_{2})$
$=2b_{2}[(b_{2}-\beta_{2\cdot 1})var(X_{2})-\beta_{1\cdot 2}cov(X_{1},X_{2})]+var(Y)$
@@
$\rho(Y^{*},X_{1})$ $=\frac{cov(Y^{*},X_{1})}{\sqrt{var(Y^{*})var(X_{1})}}$
$\approx\frac{b_{2}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{1})var(X_{2})}}{\sqrt{var(Y^{*})var(X_{1})}}$
$=\frac{b_{2}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{2})}}{\sqrt{var(Y^{*})}}$
$=\frac{b_{2}\cdot\rho(X_{1},X_{2})\sqrt{var(X_{2})}}{\sqrt{2b_{2}[(b_{2}-\beta_{2\cdot 1})var(X_{2})-\beta_{1\cdot 2}cov(X_{1},X_{2})]+var(Y)}}.$ (13)